OpenStudy (anonymous):

Complete the square for the equation: x^2+2x-3=0 a. (x+1)^2=4 b. (x+1)^2=9 c. (x+2)^2=4 d. (x+1.5)=4.25

OpenStudy (anonymous):

So..\[x^2+2x=3\]\[x^2+2x+1=3+1\]\[(x+1)^2=4\]

OpenStudy (anonymous):

Let me know if you don't understand what I did :).

OpenStudy (cwrw238):

x^2+2x-3=0 take half of coefficient of x - that is 1: (x + 1)^2 - 1 - 3 = 0 (x +1)^2 = 4 take square root of both sides to solve for x